OpenStudy (anonymous):

−4(x + 1) − 3 = −3(x − 4)

OpenStudy (anonymous):

-4x -4 -3 = -3x +12 -4x -7 = 3x +12

OpenStudy (anonymous):

-7x =12+7 -7x = 19 x=-19/7

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

I think you forgot the - in front of 3x. -4x -7 = -3x +12 -x = 5 x = 5/-1

OpenStudy (anonymous):

@xaadi1993

OpenStudy (anonymous):

u are right @jazy

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

yes